# Jets are the most robust observable ingredient of common envelope evolution

### v1

@@ 4 Summary
Most of the tens of CE simulations (e.g., just from the last decade, Nandez et al. 2014; Staff et al. 2016; Kuruwita et al. 2016; Ohlmann et al. 2016; Iaconi et al. 2017; Chamandy et al. 2019; Law-Smith et al. 2020; Glanz & Perets 2021a, b; González-Bolívar et al. 2022, 2024b; Lau et al. 2022a, b; Bermúdez-Bustamante et al. 2024b; Chamandy et al. 2024; Gagnier & Pejcha 2024; Landri et al. 2024; Rosselli-Calderon et al. 2024; Vetter et al. 2024) do not include the role of jets. Those simulations that do include jets cover a small fraction of the orbital period or omit other processes (e.g., Moreno Méndez et al. 2017; Shiber & Soker 2018; López-Cámara et al. 2019; Shiber et al. 2019; López-Cámara et al. 2020; Hillel et al. 2022, 2023; López-Cámara et al. 2022; Zou et al. 2022; Soker 2022; Schreier et al. 2023; Gurjar et al. 2024; Shiber & Iaconi 2024). I list many (but not all) of the last decade’s simulations to emphasize that despite the considerable progress and many studies in the last decade, we are still far from correctly simulating the CEE.
@@ Acknowledgements
I heavily used the catalog of PNe with a binary central star built and maintained by David Jones (see text) and the Planetary Nebula Image Catalogue (PNIC) of Bruce Balick (Balick, 2006) https://faculty.washington.edu/balick/PNIC/ A grant from the Pazy Foundation supported this research.

### v2

@@ 4 Summary
Most of the tens of CEE simulations (e.g., just from the last decade, Nandez et al. 2014; Staff et al. 2016; Kuruwita et al. 2016; Ohlmann et al. 2016; Iaconi et al. 2017; Chamandy et al. 2019; Law-Smith et al. 2020; Glanz & Perets 2021a, b; González-Bolívar et al. 2022, 2024b; Lau et al. 2022a, b; Bermúdez-Bustamante et al. 2024b; Chamandy et al. 2024; Gagnier & Pejcha 2024; Landri et al. 2024; Rosselli-Calderon et al. 2024; Vetter et al. 2024) do not include the role of jets. Those simulations that do include jets cover a small fraction of the orbital period or omit other processes (e.g., Moreno Méndez et al. 2017; Shiber & Soker 2018; López-Cámara et al. 2019; Shiber et al. 2019; López-Cámara et al. 2020; Hillel et al. 2022, 2023; López-Cámara et al. 2022; Zou et al. 2022; Soker 2022; Schreier et al. 2023; Gurjar et al. 2024; Shiber & Iaconi 2024). I list many (but not all) of the last decade’s simulations to emphasize that despite the considerable progress and many studies in the last decade, we are still far from correctly simulating the CEE.
@@ Acknowledgements
I thank Gunter Cibis and an anonymous referee for their helpful comments. I heavily used the catalog of PNe with a binary central star built and maintained by David Jones (see text) and the Planetary Nebula Image Catalogue (PNIC) of Bruce Balick (Balick 2006; https://faculty.washington.edu/balick/PNIC/). A grant from the Pazy Foundation supported this research.